question 1:joe and janna leave home at the same time traveling in opposite directions. joe drives 45 miles per hour and janna drives 40 miles per hour. in how many hours will they be 510 miles apart?

Answer by richwmiller(17219)

One question per submission please unless the others depend on the previous.
But you even comment that they are different topics.
Keep to the same topic.
both distances will total 510
the times will be the same
45x+40x=510
combine
85x=510
divide by 5 just to make the numbers smaller (not a necessary step)
17x=102
x=102/17=6 hours
45*6+40*6=510
270+240=510
at 45 mph joe went 270 miles and janna went 240 miles
after 6 hours they were 510 miles apart.